# Covariate-Adaptive Randomization in Clinical Trials without Inflated Variances

### v2

@@ 1 Introduction
Covariate adaptive randomization (CAR) procedures have often been implemented for balancing treatment allocation over influential covariates in clinical trials (McEntegart, , 2003; Rosenberger and Lachin, , 2015; Taves, , 2010; Lin, Zhu and Su, , 2015). The main purposes are to enhance the comparability of treatment groups and to increase the efficiency of the statistical inference. Many CAR procedures have been proposed to balance treatment assignments within strata and over covariate margins, provided that the covariates under consideration are discrete (categorical with two or more levels) (Pocock and Simon, , 1975; Antognini and Zagoraiou, , 2011; Hu and Hu, , 2012; Rosenberger and Lachin, , 2015). The theories on CAR randomization procedures are established progressively in the past decade(Hu and Hu, , 2012; Ma, Hu and Zhang, , 2015; Hu and Zhang, , 2020; Hu, Ye and Zhang, , 2023). For balancing continuous covariates, besides the method of discretization, a variety of methods have been proposed by minimization of prescribed imbalance measures (Hoehler, , 1987; Nishi and Takaichi, , 2004; Endo et al., , 2006; Stigsby and Taves, , 2010; Ciolino et al., , 2011; Lin and Su, , 2012; Ma and Hu, , 2013; Qin et al., , 2018; Jiang, Ma and Yin, , 2018; Li, Zhou and Hu, , 2019). Recently, Ma et al, (2024) unified a lot of CAR procedures by proposing a general CAR procedure for balancing specified covariate features $\bm{\phi}(\bm{X}_{i})$ between two treatments $1$ and $2$, where $\bm{\phi}(\bm{X}_{i})=(\phi_{1}(\bm{X}_{i}),\ldots,\phi_{q}(\bm{X}_{i}))$ can contain additional covariate features besides the covariate vector $\bm{X}_{i}=(X_{i,1},\ldots,X_{i,p})$ of a unit in clinical trals, such as quadratic and interaction terms. It is shown that the convergence rate of the imbalance vectors $\sum_{i=1}^{n}(T_{i}-1/2)\bm{\phi}(\bm{X}_{i})$ is $o_{P}(n^{\epsilon_{0}})$, $\epsilon_{0}<1/2$, where $n$ is the sample size, $\bm{X}_{i}$ is the covariate vector of the $i$-th unit, and $T_{i}=1$, 0 stands for the $i$-unit being allocated to treatment $1$ and $2$, respectively. Under some suitable conditions on the density of $\bm{\phi}(\bm{X}_{i})$, $\sum_{i=1}^{n}(T_{i}-1/2)\bm{\phi}(\bm{X}_{i})$ can attain the fastest convergence rate $O_{P}(1)$, and at the same time, for any unspecified covariate feature $m(\bm{X}_{i})$, the asymptotic normality of its normalized imbalance $\sum_{i=1}^{n}(T_{i}-1/2)m(\bm{X}_{i})/\sqrt{n}$ holds and the asymptotic variance $\sigma_{m}^{2}$ is nonzero when $m(\bm{X}_{i})$ is not a linear function of $\bm{\phi}(\bm{X}_{i})$. However, $\sigma_{m}^{2}$ may exceed $\textsf{E}[m^{2}(\bm{X}_{i})]/4$ that is the variance under the complete randomization in which each unit is allocated to treatments $1$ and $2$ with equal probability $1/2$. This issue is referred to as the variance inflation. The variance inflation under the procedure of Pocock and Simon, (1975) is also founded by Zhao et al., (2024). Because the asymptotic variance of the test of treatment effects is usually a function of $\sigma_{m}^{2}$ for some unspecified covariate feature $m(\bm{X}_{i})$, the variance inflation causes such test being not valid (Ma et al, , 2024; Zhang, , 2023). The closed form of $\sigma_{m}^{2}$ is also unknown so that estimating it and adjusting the test are not easy works. A lot of methods for the adjustment to obtain robust statistical inferences in the CAR clinical trials have being proposed(Bugni, Canay and Shaikh, , 2018; Ye and Shao, , 2020; Liu and Yang, , 2020; Ma, Tu and Liu, , 2024; Wang et al., , 2023). The purpose of this paper is to propose a new kind of CAR procedures so that the specified covariate features $\bm{\phi}(\bm{X}_{i})$ can be balanced well and at the same time the asymptotic variance $\sigma_{m}^{2}$ of the normalized imbalance of any unspecified covariate feature $m(\bm{X}_{i})$ is not inflated and has a closed form.
@@
From the proofs (Hu and Zhang, , 2020; Ma et al, , 2024; Zhang, , 2023), we can find that the convergence of $\sum_{i=1}^{n}(T_{i}-1/2)m(\bm{X}_{i})/\sqrt{n}$ and the finiteness of its asymptotic variance $\sigma_{m}^{2}$ are strongly based on the theory of Markov chains and the property that the allocation probabilities are symmetric about treatments so that, as a Markov chain, $\sum_{i=1}^{n}(T_{i}-1/2)\bm{\phi}(\bm{X}_{i})$ has a symmetric invariant distribution and thus $(T_{i}-1/2)m(\bm{X}_{i})$s have zero means under the invariance distribution for any $m(\bm{X}_{i})$. Recently, Liu, Hu and Ma, (2025) generalized Ma et al, (2024)’s procedure to balance covariances between two treatments with a ratio $\rho:(1-\rho)$ instead of half to half. The convergence rate of the imbalance vectors $\sum_{i=1}^{n}(T_{i}-\rho)\bm{\phi}(\bm{X}_{i})$ of the specified covariate feature $\bm{\phi}(\bm{X}_{i})$ is also shown to be $o_{P}(n^{\epsilon_{0}})$, $\epsilon_{0}<1/2$. However, by their theory, for an unspecified covariate feature $m(\bm{X})$, $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})/n$ may converge to a nonzero constant $c_{m}$ due to the fact that $\sum_{i=1}^{n}(T_{i}-\rho)\bm{\phi}(\bm{X}_{i})$ is no longer a symmetric Markov chain when $\rho\neq 1/2$. This issue is refer to as ”shift problem” of covariate $m(\bm{X}_{i})$. The shift problem is a bad news because in testing hypotheses in the covaraite-adaptive randomized clinical trials, $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})=o_{P}(n)$ is a basic condition (Ma, Hu and Zhang, , 2015; Ma et al., , 2020; Ma et al, , 2024; Zhu et al, , 2025; Liu and Hu, , 2023; Liu, Hu and Ma, , 2025), and sometimes the convergence rate $O_{P}(\sqrt{n})$ is needed (Ma, Tu and Liu, , 2024). We will show that our proposed CAR procedures are valid for all the case of $0<\rho<1$. Under the proposed CAR procedure, the ”shift problem” will no longer appear and the asymptotic variance of $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})/\sqrt{n}$ will be always controlled by $\rho(1-\rho)\textsf{E}[m^{2}(\bm{X}_{i})]$, the variance under the simple randomization in which each unit is randomized to treatments $1$ and $2$ respectively with $\rho$ and $1-\rho$, for any $m(\bm{X}_{i})$.
@@
In the next section, the framework of the new CAR procedures are proposed with theories on its properties. In Section 3, as an application, the properties of a simple test of treatment effects under the proposed CAR procedure are studied based on the observed responses on treatments. It is shown that the test always controls the type I error rate without any information of the covariates, and the test can be adjusted to have the precise the type I error rate and increase the power when the data of the covariate features $\bm{\phi}(\bm{X}_{i})$s in the randomization stage are also available in the analysis stage. The proofs are given in the last section.
@@ 2.1 Framework
Suppose that $n$ experimental units are to be assigned to two treatment groups. Let $T_{i}$ be the assignment of the $i$-th unit, i.e., $T_{i}=1$ for treatment 1 and $T_{i}=0$ for treatment 0. Consider $p$ covariates available for each unit and use $\bm{X}_{i}=(X_{i,1},...,X_{i,p})\in\mathbb{R}^{p}$ to denote the covariates of the $i$-th unit. Besides $\bm{X}_{i}$, the unit may have other random features such as responses and other covaraites which are observable or unobservable and not considered in the randomization, we denote them by $Z_{i},W_{i}$. We assume that $\{(\bm{X}_{i},Z_{i},W_{i});i=1,2,\ldots,n\}$ are independent and identically distributed as a random vector $(\bm{X},W,Z)$. We consider a procedure to balance general covariate features $\bm{\phi}(\bm{X}_{i})$, defined by a feature map $\bm{\phi}(\bm{x}):\mathbb{R}^{p}\to\mathbb{R}^{q}$ that maps $\bm{X}_{i}$ into a $q$-dimensional feature space. Here, $q$ is usually larger than $p$ so that $\bm{\phi}(\bm{X}_{i})$ has more features than the original covariates. The vector $\bm{\Lambda}_{n}=\sum_{i=1}^{n}(T_{i}-\rho)\bm{\phi}(\bm{X}_{i})$ is the imbalance measure of the covariate feature $\bm{\phi}(\bm{X})$, where $0<\rho<1$. The numerical imbalance measure $Imb_{n}$ is defined as the squared Euclidean norm of the imbalance vector $\bm{\Lambda}_{n}$, i.e.,
@@
$=$ $2(T_{n}-\rho)\langle\bm{\Lambda},\phi(\bm{X}_{n-1})\rangle+(T_{n}-\rho)^{2}\|\bm{\phi}(\bm{X}_{n})\|^{2}.$
@@
According to the theoretical results given in Theorem 2.2 and Remark 2.2, the parameter $\gamma$ controls the size of the imbalance measure $Imb_{n}$. The smaller the parameter $\gamma$ , the smaller is $Imb_{n}$, but possibly the larger is the imbalance $\|\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}\|^{2}$ of the unspecified feature $Z_{i}$ that is not a linear fuction of $\bm{\phi}(\bm{X}_{i})$. The value of $\gamma$ is suggested to be in $[0.5,1)$ so that the asymptotic expectation of $\|\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}\|^{2}/n$ does not exceed the one under the simple randomization for any $Z_{i}$s.
@@ 2.2 Theoretical Properties
In this subsection we shall study the theoretical properties of the CAR procedures with a feature map $\bm{\phi}(\bm{x})=(\phi_{1}(\bm{x}),\ldots,\phi_{q}(\bm{x})):\mathbb{R}^{p}\mapsto\mathbb{R}^{q}$. Our goal is to obtain the convergence rate of covariate imbalance measured by $\bm{\Lambda}_{n}=\sum_{i=1}^{n}(2T_{i}-1)\bm{\phi}(\bm{X}_{i})$ and the asymptotic normality of $\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}$.
@@
The sequence $\{(\bm{X}_{i},Z_{i},W_{i})\}$ is a sequence of independent and identically distributed as a random vector $(\bm{X},W,Z)$..
@@
In the sequel of this paper, for sequences of real numbers $\{a_{n}\}$, $\{b_{n}\}$ $(b_{n}>0)$ and a sequence of random vectors $\{\bm{Y}_{n}\}$, $a_{n}=O(b_{n})$ means $|a_{n}/b_{n}|\leq C$ for some $C>0$, $a_{n}=o(b_{n})$ means $a_{n}/b_{n}\to 0$, $\bm{Y}_{n}=O(b_{n})$ (resp. $o(b_{n})$) in $L_{r}$ means $(\textsf{E}\|\bm{Y}_{n}\|^{r})^{1/r}=O(b_{n})$ (resp. $o(b_{n})$), $\bm{Y}_{n}=o_{P}(b_{n})$ means that $\|\bm{Y}_{n}\|/b_{n}\to 0$ in probability, and $\bm{Y}_{n}=O_{P}(b_{n})$ means that $\|\bm{Y}_{n}\|/b_{n}$ is bounded in probability.
@@
Suppose that Assumptions 2.1 and 2.2 are satisfied. Then
$\textsf{E}[\|\bm{\Lambda}_{n}\|^{2}]=O\big(n^{\gamma}+n^{\frac{1}{2}+\frac{1}{2(r_{0}-1)}}\big)=o(n).$ (2.5)
If $\textsf{E}[|Z|]<\infty$ and $\gamma>\frac{1}{4}+\frac{1}{4(r_{0}-1)}$, then
@@
Moreover, if $r_{0}\geq 4$, $\gamma>\frac{1}{2}+\frac{1}{4(r_{0}-1)}$, $\textsf{E}Z^{2}<\infty$, $\textsf{E}W^{2}<\infty$ and $\textsf{E}[W]=0$, then
@@
where $\vec{\sigma}_{Z}^{2}=(1-\rho)\rho\textsf{E}[\widetilde{Z}^{2}]$, $\widetilde{Z}=Z-\langle\bm{\phi}(\bm{X}),\bm{x}_{0}\rangle$ and $\bm{x}_{0}=\arg\min\{\textsf{E}\big[(Z-\langle\bm{\phi}(\bm{X}),\bm{x}\rangle)^{2}\big]:\bm{x}\in\mathbb{R}^{q}\}$.
@@
Liu, Hu and Ma, (2025) proposes a CAR procedure for balancing covariate features $\bm{\phi}(\bm{X}_{i})$ between treatments with an unequal ratio $\rho:(1-\rho)$. When the additional covariate $m(\bm{X}_{i})$ is not the linear function of $\bm{\phi}(\bm{X}_{i})$, it may happen that $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})\overset{P}{\to}c\neq 0$. This issue is referred as ”shift problem”. The simulation results presented in the Supplementary Material of Liu, Hu and Ma, (2025) show that the shift problem on the additional covariates is common when balancing continuous covariates by using their CAR procedure. Now, (2.7) shows that the shift problem no longer exists under our proposed CAR procedure.
@@
The simple randomization is randomizing each patient to treatment $1$ with the same probability $\rho$. For the simple randomization, $\vec{\sigma}_{Z}^{2}=\rho(1-\rho)\textsf{E}Z^{2}$. So, the proposed CAR procedure always has smaller asymptotic variance than the simple randomization for all observed or unobserved feature $Z_{i}$ with $\textsf{Pj}[Z|\bm{\phi}(\bm{X})]\neq 0$. The variance inflation is no longer a problem under our proposed CAR procedure.
@@
The normalized imbalance $\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}/\sqrt{n}$ of $Z_{i}$s is an endogenous variable of the design, and the normalized sum $\sum_{i=1}^{n}W_{i}/\sqrt{n}$ of $W_{i}$s can be regraded as an exogenous variable. (2.8) means that they are asymptotic independent regardless $W_{i}$ and $Z_{i}$ themselves are independent or not.
When $\bm{\phi}(\bm{X})$ has finite moments of all orders. The convergence rate of the imbalance $\bm{\Lambda}_{n}$ can be improved a lot.

### v3

@@ 1 Introduction
Covariate adaptive randomization (CAR) procedures have often been implemented for balancing treatment allocation over influential covariates in clinical trials (McEntegart, , 2003; Rosenberger and Lachin, , 2015; Taves, , 2010; Lin, Zhu and Su, , 2015). The main purposes are to enhance the comparability of treatment groups and to increase the efficiency of the statistical inference. Many CAR procedures have been proposed to balance treatment assignments within strata and over covariate margins, provided that the covariates under consideration are discrete (categorical with two or more levels) (Pocock and Simon, , 1975; Antognini and Zagoraiou, , 2011; Hu and Hu, , 2012; Rosenberger and Lachin, , 2015). The theories on CAR randomization procedures are established progressively in the past decade(Hu and Hu, , 2012; Ma, Hu and Zhang, , 2015; Hu and Zhang, , 2020; Hu, Ye and Zhang, , 2023). For balancing continuous covariates, besides the method of discretization, a variety of methods have been proposed by minimization of prescribed imbalance measures (Hoehler, , 1987; Nishi and Takaichi, , 2004; Endo et al., , 2006; Stigsby and Taves, , 2010; Ciolino et al., , 2011; Lin and Su, , 2012; Ma and Hu, , 2013; Qin et al., , 2018; Jiang, Ma and Yin, , 2018; Li, Zhou and Hu, , 2019). Recently, Ma et al, (2024) unified a lot of CAR procedures by proposing a general CAR procedure for balancing specified covariate features $\bm{\phi}(\bm{X}_{i})$ between two treatments $1$ and $2$, where $\bm{\phi}(\bm{X}_{i})=(\phi_{1}(\bm{X}_{i}),\ldots,\phi_{q}(\bm{X}_{i}))$ can contain additional covariate features besides the covariate vector $\bm{X}_{i}=(X_{i,1},\ldots,X_{i,p})$ of a unit in clinical trals, such as quadratic and interaction terms. It is shown that the convergence rate of the imbalance vectors $\sum_{i=1}^{n}(T_{i}-1/2)\bm{\phi}(\bm{X}_{i})$ is $o_{P}(n^{\epsilon_{0}})$, $\epsilon_{0}<1/2$, where $n$ is the sample size, $\bm{X}_{i}$ is the covariate vector of the $i$-th unit, and $T_{i}=1$, 0 stands for the $i$-unit being allocated to treatment $1$ and $2$, respectively. Under some suitable conditions on the density of $\bm{\phi}(\bm{X}_{i})$, $\sum_{i=1}^{n}(T_{i}-1/2)\bm{\phi}(\bm{X}_{i})$ can attain the fastest convergence rate $O_{P}(1)$, and at the same time, for any unspecified covariate feature $m(\bm{X}_{i})$, the asymptotic normality of its normalized imbalance $\sum_{i=1}^{n}(T_{i}-1/2)m(\bm{X}_{i})/\sqrt{n}$ holds and the asymptotic variance $\sigma_{m}^{2}$ is nonzero when $m(\bm{X}_{i})$ is not a linear function of $\bm{\phi}(\bm{X}_{i})$. However, $\sigma_{m}^{2}$ may exceed $\textsf{E}[m^{2}(\bm{X}_{i})]/4$ that is the variance under the complete randomization in which each unit is allocated to treatments $1$ and $2$ with equal probability $1/2$. This issue is referred to as the variance inflation. The variance inflation under the procedure of Pocock and Simon, (1975) is found by Zhao et al., (2024). Because the asymptotic variance of the test of treatment effects is usually a function of $\sigma_{m}^{2}$ for some unspecified covariate feature $m(\bm{X}_{i})$, the variance inflation causes such test being not valid (Ma et al, , 2024; Zhang, , 2023). The closed form of $\sigma_{m}^{2}$ is also unknown so that estimating it and adjusting the test are not easy works. A lot of methods for the adjustment to obtain robust statistical inferences in the CAR clinical trials have being proposed(Bugni, Canay and Shaikh, , 2018; Ye and Shao, , 2020; Liu and Yang, , 2020; Ma, Tu and Liu, , 2024; Wang et al., , 2023). The purpose of this paper is to propose a new kind of CAR procedures so that the specified covariate features $\bm{\phi}(\bm{X}_{i})$ can be balanced well and at the same time the asymptotic variance $\sigma_{m}^{2}$ of the normalized imbalance of any unspecified covariate feature $m(\bm{X}_{i})$ is not inflated and has a closed form.
@@
From the proofs (Hu and Zhang, , 2020; Ma et al, , 2024; Zhang, , 2023), we can find that the convergence of $\sum_{i=1}^{n}(T_{i}-1/2)m(\bm{X}_{i})/\sqrt{n}$ and the finiteness of its asymptotic variance $\sigma_{m}^{2}$ are strongly based on the theory of Markov chains and the property that the allocation probabilities are symmetric about treatments so that, as a Markov chain, $\sum_{i=1}^{n}(T_{i}-1/2)\bm{\phi}(\bm{X}_{i})$ has a symmetric invariant distribution and thus $(T_{i}-1/2)m(\bm{X}_{i})$s have zero means under the invariance distribution for any $m(\bm{X}_{i})$. Recently, Liu, Hu and Ma, (2025) generalized Ma et al, (2024)’s procedure to balance covariances between two treatments with a ratio $\rho:(1-\rho)$ instead of half to half. The convergence rate of the imbalance vectors $\sum_{i=1}^{n}(T_{i}-\rho)\bm{\phi}(\bm{X}_{i})$ of the specified covariate feature $\bm{\phi}(\bm{X}_{i})$ is also shown to be $o_{P}(n^{\epsilon_{0}})$, $\epsilon_{0}<1/2$. However, by their theory, for an unspecified covariate feature $m(\bm{X})$, $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})/n$ may converge to a nonzero constant $c_{m}$ due to the fact that $\sum_{i=1}^{n}(T_{i}-\rho)\bm{\phi}(\bm{X}_{i})$ is no longer a symmetric Markov chain when $\rho\neq 1/2$. This issue is refer to as ”shift problem” of covariate $m(\bm{X}_{i})$. The shift problem is a bad news because in testing hypotheses in the covaraite-adaptive randomized clinical trials, $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})=o_{P}(n)$ is a basic condition (Ma, Hu and Zhang, , 2015; Ma et al., , 2020; Ma et al, , 2024; Zhu et al, , 2025; Liu and Hu, , 2023; Liu, Hu and Ma, , 2025), and sometimes the convergence rate $O_{P}(\sqrt{n})$ is needed (Ma, Tu and Liu, , 2024). We will show that our proposed CAR procedures are valid for all the case of $0<\rho<1$. Under the proposed CAR procedure, the ”shift problem” will no longer appear and the asymptotic variance of $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})/\sqrt{n}$ will be always controlled by $\rho(1-\rho)\textsf{E}[m^{2}(\bm{X}_{i})]$, the variance under the simple randomization in which each unit is randomized to treatments $1$ and $2$ respectively with probabilities $\rho$ and $1-\rho$, for any $m(\bm{X}_{i})$.
@@
In the next section, the framework of the new CAR procedures are proposed with theories on its properties. In Section 3, as an application, the properties of a simple test of treatment effects under the proposed CAR procedure are studied based on the observed responses on treatments. It is shown that the test always controls the type I error rate without any information of the covariates, and the test can be adjusted to have the precise type I error rate and increase the power when the data of the covariate features $\bm{\phi}(\bm{X}_{i})$s in the randomization stage are also available in the analysis stage. The proofs are given in the last section.
@@ 2.1 Framework
Suppose that $n$ experimental units are to be assigned to two treatment groups. Let $T_{i}$ be the assignment of the $i$-th unit, i.e., $T_{i}=1$ for treatment 1 and $T_{i}=0$ for treatment 0. Consider $p$ covariates available for each unit and use $\bm{X}_{i}=(X_{i,1},...,X_{i,p})\in\mathbb{R}^{p}$ to denote the covariates of the $i$-th unit. Besides $\bm{X}_{i}$, the unit may have other random features such as responses and other covaraites which are observable or unobservable and not considered in the randomization, we denote them by $Z_{i},W_{i}$. We assume that $\{(\bm{X}_{i},Z_{i},W_{i});i=1,2,\ldots,n\}$ are independent and identically distributed as a random vector $(\bm{X},W,Z)$. We consider a procedure to balance general covariate features $\bm{\phi}(\bm{X}_{i})$, defined by a feature map $\bm{\phi}(\bm{x}):\mathbb{R}^{p}\to\mathbb{R}^{q}$ that maps $\bm{X}_{i}$ into a $q$-dimensional feature space. Here, $q$ is usually larger than $p$ so that $\bm{\phi}(\bm{X}_{i})$ has more features than the original covariates. The purpose of the covariate-adaptive randomization is to balance $\bm{\phi}(\bm{X}_{i})$s between treatments $1$ and $2$ with a ratio $\rho:(1-\rho)$, where $0<\rho<1$. The vector $\bm{\Lambda}_{n}=\sum_{i=1}^{n}(T_{i}-\rho)\bm{\phi}(\bm{X}_{i})$ is the imbalance measure of the covariate feature $\bm{\phi}(\bm{X})$. The numerical imbalance measure $Imb_{n}$ is defined as the squared Euclidean norm of the imbalance vector $\bm{\Lambda}_{n}$, i.e.,
@@
$=$ $2(T_{n}-\rho)\langle\bm{\Lambda}_{n-1},\phi(\bm{X}_{n})\rangle+(T_{n}-\rho)^{2}\|\bm{\phi}(\bm{X}_{n})\|^{2}.$
@@
According to the theoretical results given in Theorems 2.1 and 2.3, the parameter $\gamma$ controls the size of the imbalance measure $Imb_{n}$. The smaller the parameter $\gamma$ , the smaller is $Imb_{n}$, but possibly the larger is the imbalance $\|\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}\|^{2}$ of the unspecified feature $Z_{i}$ that is not a linear fuction of $\bm{\phi}(\bm{X}_{i})$. The value of $\gamma$ is suggested to be in $[0.5,1)$ so that the asymptotic expectation of $\|\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}\|^{2}/n$ does not exceed the one under the simple randomization for any $Z_{i}$s.
@@ 2.2 Theoretical Properties
In this subsection we shall study the theoretical properties of the CAR procedures with a feature map $\bm{\phi}(\bm{x})=(\phi_{1}(\bm{x}),\ldots,\phi_{q}(\bm{x})):\mathbb{R}^{p}\mapsto\mathbb{R}^{q}$. Our goal is to obtain the convergence rate of covariate imbalance measured by $\bm{\Lambda}_{n}=\sum_{i=1}^{n}(T_{i}-\rho)\bm{\phi}(\bm{X}_{i})$ and the asymptotic normality of $\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}$.
@@
The sequence $\{(\bm{X}_{i},Z_{i},W_{i}),i=1,2,\ldots\}$ is a sequence of independent and identically distributed as a random vector $(\bm{X},W,Z)$.
@@
In the sequel of this paper, for sequences of real numbers $\{a_{n}\}$, $\{b_{n}\}$ $(b_{n}>0)$ and a sequence of random vectors $\{\bm{Y}_{n}\}$, $a_{n}=O(b_{n})$ means $|a_{n}/b_{n}|\leq C$ for some $C>0$, $a_{n}=o(b_{n})$ means $a_{n}/b_{n}\to 0$, $\bm{Y}_{n}=O(b_{n})$ (resp. $o(b_{n})$) in $L_{r}$ means $(\textsf{E}\|\bm{Y}_{n}\|^{r})^{1/r}=O(b_{n})$ (resp. $o(b_{n})$), $\bm{Y}_{n}=o_{P}(b_{n})$ means that $\|\bm{Y}_{n}\|/b_{n}\to 0$ in probability, and $\bm{Y}_{n}=O_{P}(b_{n})$ means that $\|\bm{Y}_{n}\|/b_{n}$ is bounded in probability. A positive constant which depends on $\alpha$ and $\beta$ is denoted by $C_{\alpha,\beta}$.
@@
Suppose that Assumptions 2.1 and 2.2 are satisfied.
We have
$\left(\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r}\right]\right)^{1/r}\leq C_{r,r_{0}}\Big(n^{\frac{\gamma}{2}}+n^{\frac{\gamma}{2}-\frac{(r_{0}+2-r)\gamma-2}{2r(r_{0}+1-r)}\big)}\Big),\;2\leq r\leq r_{0},$ (2.5)
$\left(\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r}\right]\right)^{1/r}\leq C_{r,r_{0}}\Big(n^{\frac{\gamma}{r}}+n^{\frac{(r_{0}-2)\gamma+2}{2r(r_{0}+1-r)}}\Big),\;\;1\leq r\leq 2.$ (2.6)
In particular,
$\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{2}\right]=O\big(n^{\gamma}+n^{\frac{(r_{0}-2)\gamma+2}{2(r_{0}-1)}}\big)\text{ and }\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r_{0}}\right]=O(n^{\frac{\gamma r_{0}}{2}+1-\gamma})=o(n^{r_{0}/2}).$ (2.7)
If $\textsf{E}[|Z|]<\infty$ and $\gamma>\frac{2}{3r_{0}-2}$ (more precisely, $\gamma>\frac{4}{r_{0}^{2}+5}$ when $2<r_{0}\leq 3$, $\gamma>\frac{2}{3r_{0}-2}$ when $3\leq r_{0}\leq 4$, $\gamma>\frac{8}{(r_{0}+2)^{2}+4}$ when $r_{0}\geq 4$), then
@@
If $r_{0}\geq 4$, $\gamma\geq 2/r_{0}$ and $\textsf{E}Z^{2}<\infty$, then
$\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}=$ $\sum_{i=1}^{n}\big((T_{i}-\rho)\widetilde{Z}_{i}-\textsf{E}[(\ell_{i}-\rho)\widetilde{Z}_{i}|\bm{\Lambda}_{i-1}]\big)$
$+O(n^{\gamma/2}+n^{1-\gamma})\text{ in }L_{1},$ (2.10)
where $\widetilde{Z}=Z-\langle\bm{\phi}(\bm{X}),\bm{x}_{0}\rangle$ and $\bm{x}_{0}=\arg\min\{\textsf{E}\big(Z-\langle\bm{\phi}(\bm{X}),\bm{x}\rangle\big)^{2}:\bm{x}\in\mathbb{R}^{q}\}$.
If $r_{0}\geq 4$, $\gamma>\frac{1}{2}$, $\textsf{E}Z^{2}<\infty$, $\textsf{E}W^{2}<\infty$ and $\textsf{E}[W]=0$, then
@@
where $\vec{\sigma}_{Z}^{2}=\rho(1-\rho)\textsf{E}[\widetilde{Z}^{2}]$, and
when $(r_{0}-2)\gamma\geq 2$.
@@
Liu, Hu and Ma, (2025) proposes a CAR procedure for balancing covariate features $\bm{\phi}(\bm{X}_{i})$ between treatments with an unequal ratio $\rho:(1-\rho)$. When the additional covariate $m(\bm{X}_{i})$ is not the linear function of $\bm{\phi}(\bm{X}_{i})$, there is no theoretical guarantee for $\frac{1}{n}\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})\overset{P}{\to}0$, it may happen that $\frac{1}{n}\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})\overset{P}{\to}c\neq 0$. This issue is referred as ”shift problem”. The simulation results presented in the Supplementary Material of Liu, Hu and Ma, (2025) show that the shift problem on the additional covariates is common when balancing continuous covariates by using their CAR procedure. Now, (2.9) shows that the shift problem no longer exists under our proposed CAR procedure.
@@
The simple randomization is a randomization procedure which randomizes each unit to treatment $1$ with the same probability $\rho$. For the simple randomization, $\vec{\sigma}_{Z}^{2}=\rho(1-\rho)\textsf{E}Z^{2}$. So, the proposed CAR procedure always has smaller asymptotic variance than the simple randomization for all observed or unobserved feature $Z_{i}$ with $\textsf{Pj}[Z|\bm{\phi}(\bm{X})]\neq 0$. The variance inflation is no longer a problem under our proposed CAR procedure.
@@
The normalized imbalance $\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}/\sqrt{n}$ of $Z_{i}$s is an endogenous variable of the design, and the normalized sum $\sum_{i=1}^{n}W_{i}/\sqrt{n}$ of $W_{i}$s can be regraded as an exogenous variable. (2.11) means that they are asymptotically independent regardless $W_{i}$ and $Z_{i}$ themselves are independent or not.
Remark 2.2.
If $\ell(x)$ is thrice differentiable at $x=0$, and $\ell^{\prime\prime}(0)=0$ or $\textsf{E}[(\bm{\phi}(\bm{X}))^{\prime}\bm{\phi}(\bm{X})\widetilde{Z}]=\bm{0}$, then (2.11) remains true when $r_{0}\geq 7$ and $1/3<\gamma<1$.
If $\textsf{E}[Z|\bm{\phi}(\bm{X})]\in Span\{\bm{\phi}(\bm{X})\}$, then $\textsf{E}[\widetilde{Z}|\bm{\phi}(\bm{X})]=0$ and thus $\textsf{E}[(\bm{\phi}(\bm{X}))^{\prime}\bm{\phi}(\bm{X})\widetilde{Z}]=\bm{0}$.
Next, we propose an allocation function such that (2.11) may hold for all $0<\gamma<1$. The function is defined by
where $\lambda>0$ is a tuning parameter. In (2.13), $\Phi(-x+u_{\rho/2})$, $\Phi(-x+u_{(\rho+1)/2})$ can be respectively replaced by $\alpha$, $\beta$ with $0<\alpha<\rho<\beta<1$.
Theorem 2.2.
Suppose Assumption 2.1, and that 2.2 is satisfied with $r_{0}\geq 4$, $\textsf{E}Z^{2}<\infty$, $\textsf{E}W^{2}<\infty$ and $\textsf{E}W=0$. Let $\ell(x)$ be defined as in (2.13). Then (2.11) and (2.12) hold if $(r_{0}+2)\gamma>4$.